OpenStudy (anonymous):

A woman wishes to rent a house within 9 miles of her work. If she lives x miles from her work, her transportation cost will be cx dollars per year, while her rent will be 36c/(x + 1) dollars per year, where c is a constant taking various situational factors into account. How far should she live from work to minimize her combined expenses for rent and transportation? Use the methods outlined in this section to find the minimum. x= ? mi

OpenStudy (anonymous):

Are you familiar with differential calculus? If you are, take the first derivative of y = f(x) = cx + 36c/(x + 1) .

OpenStudy (anonymous):

so would that be, c+ (-36c/(x^2 + 2x +1))

OpenStudy (anonymous):

Yes, and just set that = 0

OpenStudy (anonymous):

You will get a value for x that will contain an expression for "c".

OpenStudy (anonymous):

5c/6 and -7c/6 ?

OpenStudy (anonymous):

problem out to an intermediate step and it looks like the "c" cancels. It boils down to finding the zeros for x^2 + 2x - 35 = 0. Find the x value that gives the minimum.

OpenStudy (anonymous):

c=0?

OpenStudy (anonymous):

oh ok, it's 5! thanks so much!!

OpenStudy (anonymous):

You're welcome!